OpenStudy (anonymous):

Simplify the expression 7√7 - 2√28 A) 3 square root of 28 B)- square root 7 C) 5 square root of 28 D) 3 square root of 7

OpenStudy (anonymous):

\[2\sqrt{28} = 2\sqrt{7 * 4} = 2 * \sqrt{7} * \sqrt{4} = 2 * 2 * \sqrt{7} = 4 \sqrt{7}\] does this derivation help?

OpenStudy (anonymous):

kind of, i may need more to understand though, im a detailed learner

OpenStudy (anonymous):

me to @alexxkayee

OpenStudy (anonymous):

if you agree with the derivation above, then the problem becomes \[7 \sqrt{7} - 4 \sqrt{7}\] see how this simplifies even more?